A boy shoves his stuffed toy zebra down a frictionless chute, starting at a height of 1.91 m above the bottom of the chute and with an initial speed of 1.93 m/s. The toy animal emerges horizontally from the bottom of the chute and continues sliding along a horizontal surface with coefficient of kinetic friction 0.263. How far from the bottom of the chute does the toy zebra come to rest? Take g = 9.81 m/s2. Take into account initial kinetic energy.
Answer #1

Apply conservation of energy to solve the problem

at starting point boy has both potential energy and kinetic energy,

PEi + KEi = work done against friction

mgh + 1/2 m v^2 = fk d

mgh + 1/2 * m v^2 = uk mg d

d = mgh + 1/2 * m v^2/ uk mg

= gh + 0.5 v^2/uk g

= 9.81 (1.91) +0.5 (1.93)^2/0.263 (9.81)

=7.98 m